Vernes novel, Twenty Thousand Leagues Under the Sea, Captain Nemo and his passengers undergo many adventures as they travel the Earths oceans (a) If 1.00 league equals 3.500 km, find the depth in meters to which the crew traveled if they actually went 2.000 x 10+ leagues below the ocean surface. (b) Find the difference between your answer to part (a) and the radius of the Earth, 6.38 x 106 m. (Incidentally, author Jules Verne meant that the total distance traveled, and not the depth, was 20,000 leagues
Answer #1

a) 2*10^4 leagues= 2*10^4* 3.5*10^3= 7*10^7 m

b) required answer= 7*10^7 - 6.38*10^6 = 63.62*10^6 m

============

Comment in case any doubt .. good luck